OpenStudy (anonymous):

How do you graph this equation Y=3x-2

OpenStudy (campbell_st):

its a straight line the easiest thing to do is create a table of values pick 3 values for x, substitute them into the equation to find y then you'll have 3 ordered pairs to plot. Draw a line through all 3 points for the line so try x | 0 | 1 | 2 | -------------------- y | | | 4 | so using y = 3x - 2 and substituting x = 2 you get y = 3*2 - 2 = 4 so an ordered pair is (2, 4) you can plot this in the number plane. the reason you choose 3 points is to eliminate mistakes, the line must go through all 3 points. otherwise you need to go and check the substitutions hope it helps